OpenStudy (anonymous):

The second side of a triangle is 3 more meters than the first side. The third side is twice more than the first side. If the perimeter of a triangle is 47 meters, find the length of the second side.

OpenStudy (anonymous):

Let the first side be triangle be x Then second side of a triangle = x+3 and accordingly third side = 2x Perimeter of the triangle = sum of all the sides = 47 x+x+3+2x = 47 4x+3 = 47 Subtracting 3 from either side results in.. 4x+3-3 = 47-3 4x= 44 Dividing both sides by 4 x=11 So second side of a triangle = x+3 = 11+3 Ans : 14 meter

OpenStudy (anonymous):

please help with formula

OpenStudy (dumbcow):

there is no formula...except maybe that perimeter = sum of 3 sides